If the width of a rectangle is 1 inch less than its length and the diagonal is 1 inch longer than its length, find the length and the width.

Answer by RAY100(1637)

Let,,,Length = L
Width = L-1
Diagonal = L+1
.
.
Diagonal is hypotenuse of a right triangle, therefore use the pythagorian theorem,
.
c^2 = a^2 + b^2
.
(L+1)^2 = (L)^2 + (L-1)^2
.
L^2 +2L +1 = L^2 + L^2 -2L +1
.
L^2 - 4L =0
.
L (L-4) =0
.
L=4,,,,0 (not realistic)
.
L=4, W=3, Dia = 5
.
check,,5^2 = 3^2 + 4^2 ,,,,,ok
.